Number problem. If one-half of one integer is subtracted from three-fifths of the next consecutive integer, the difference is 3. What are the two integers?
Answer by ptaylor(1894)

Let x=one integer
x/2=1/2 of one integer
Then x+1=next consecutive integer
(3/5)(x+1)=3/5 of the next consecutive integer
Now we are told that:
(3/5)(x+1)-x/2=3 multiply each term by 10 (LCM)
6(x+1)-5x=30 get rid of parens (distributive law)
6x+6-5x=30 subtract 6 from both sides
6x+6-6-5x=30-6 collect like terms
x=24 on integer
x+1=24+1=25 next consecutive integer
CK
(3/5)*(25)-(1/2)*24=3
15-12=3
3=3

Hope this helps--ptaylor